# L4DR: LiDAR-4DRadar Fusion for Weather-Robust 3D Object Detection

### v4

@@ Abstract
LiDAR-based 3D object detection is crucial for autonomous driving. However, due to the quality deterioration of LiDAR point clouds, it suffers from performance degradation in adverse weather conditions. Fusing LiDAR with the weather-robust 4D radar sensor is expected to solve this problem; however, it faces challenges of significant differences in terms of data quality and the degree of degradation in adverse weather. To address these issues, we introduce L4DR, a weather-robust 3D object detection method that effectively achieves LiDAR and 4D Radar fusion. Our L4DR proposes Multi-Modal Encoding (MME) and Foreground-Aware Denoising (FAD) modules to reconcile sensor gaps, which is the first exploration of the complementarity of early fusion between LiDAR and 4D radar. Additionally, we design an Inter-Modal and Intra-Modal ({IM}2) parallel feature extraction backbone coupled with a Multi-Scale Gated Fusion (MSGF) module to counteract the varying degrees of sensor degradation under adverse weather conditions. Experimental evaluation on a VoD dataset with simulated fog proves that L4DR is more adaptable to changing weather conditions. It delivers a significant performance increase under different fog levels, improving the 3D mAP by up to 20.0% over the traditional LiDAR-only approach. Moreover, the results on the K-Radar dataset validate the consistent performance improvement of L4DR in real-world adverse weather conditions.
@@ Conclusion
Benefiting from the sensor weather robustness of 4D radar, the 4D Radar-based method receives little performance impact from weather. Therefore, the performance of the 4D Radar-based method under each weather can be used to determine the difference in difficulty caused by the distribution of labeled objects for different weather on the K-radar dataset. As shown in the results demonstrated by the 4D Radar-based method (RTNH, PointPillars) in Tables 8 and 9, it can be found that the performance of the 4D Radar-based method is significantly worse than that of most of the bad weather in normal weather. This result illustrates that labeled objects in bad weather scenarios are much easier to detect in the K-Radar dataset. This makes it possible that even when LiDAR is degraded in bad weather, the performance is still higher than in normal weather due to the fact that the labeled targets in the scene are easier to detect.
@@
More Fusion Details
Below we present the implementation details of the individual fusion methods compared in Table 6 of the main text, all of which are implemented on the PointPiillars baseline.
Concat.
We directly concatenate the pseudo-images of LiDAR with 4DRadar in the channel dimension after PointPillar coding.
Cross-Attn.
We used a 32-dimensional sin/cos position-encoded 4-head attention layer to calculate the Cross-Modal Pillar feature added to the 4DRadar Pillar feature from the LiDAR Pillar feature to the 4DRadar P illar feature, and also to calculate the 4DRadar Pillar feature to the LiDAR P illar feature’s Cross-Modal Pillar feature and added to the LiDAR Pillar feature.
Self-Attn.
We use a 32-dimensional sin/cos position-encoded 4-head attentional layer to compute self-attentional features on the last two BEV features of the 2D BackBone and add them to the original features.
SE Block.
We use 2x Squeeze’s SEBlock to compute SE features for each BEV feature of the 2D BackBone and add them to the original feature.
CBAM Block.
We use CBAM Block to compute SE features for each BEV feature of the 2D BackBone and add them to the original feature.

### v5

@@ Abstract
LiDAR-based 3D object detection is crucial for autonomous driving. However, due to the quality deterioration of LiDAR point clouds, it suffers from performance degradation in adverse weather conditions. Fusing LiDAR with the weather-robust 4D radar sensor is expected to solve this problem; however, it faces challenges of significant differences in terms of data quality and the degree of degradation in adverse weather. To address these issues, we introduce L4DR, a weather-robust 3D object detection method that effectively achieves LiDAR and 4D Radar fusion. Our L4DR proposes Multi-Modal Encoding (MME) and Foreground-Aware Denoising (FAD) modules to reconcile sensor gaps, which is the first exploration of the complementarity of early fusion between LiDAR and 4D radar. Additionally, we design an Inter-Modal and Intra-Modal ({IM}2) parallel feature extraction backbone coupled with a Multi-Scale Gated Fusion (MSGF) module to counteract the varying degrees of sensor degradation under adverse weather conditions. Experimental evaluation on a VoD dataset with simulated fog proves that L4DR is more adaptable to changing weather conditions. It delivers a significant performance increase under different fog levels, improving the 3D mAP by up to 20.0% over the traditional LiDAR-only approach. Moreover, the results on the K-Radar dataset validate the consistent performance improvement of L4DR in real-world adverse weather conditions. The code can be publicly available at: https://github.com/ylwhxht/L4DR.
@@ Conclusion
Benefiting from the sensor weather robustness of 4D radar, the 4D Radar-based method receives little performance impact from weather. Therefore, the performance of the 4D Radar-based method under each weather can be used to determine the difference in difficulty caused by the distribution of labeled objects for different weather on the K-radar dataset. As shown in the results demonstrated by the 4D Radar-based method (RTNH, PointPillars) in Tables 9 and 10, it can be found that the performance of the 4D Radar-based method is significantly worse than that of most of the bad weather in normal weather. This result illustrates that labeled objects in bad weather scenarios are much easier to detect in the K-Radar dataset. This makes it possible that even when LiDAR is degraded in bad weather, the performance is still higher than in normal weather due to the fact that the labeled targets in the scene are easier to detect.
@@
MME FAD {IM}2̂ MSGF Total Normal Overcast Fog Rain Sleet Light Snow Heavy Snow
37.8 39.8 46.3 59.8 28.2 31.4 50.7 24.6
51.6 46.1 58.7 77.4 52.7 43.6 63.5 42.8
51.8 49.8 60.8 77.2 50.9 43.7 64.5 33.0
52.5 51.8 63.6 66.7 51.6 46.0 52.1 36.6
53.5 53.0 64.1 73.2 53.8 46.2 52.4 37.0
Table 8: The experimental ablation results of each module on K-Radar dataset. Best in bold.
@@
We also have conducted more ablation on K-Radar dataset as in Table 8. The first row represents the RTNH as the LiDAR-only baseline for comparison. Adding MME multi-modal information significantly improves performance, especially in challenging weather. Including all modules further enhances total performance. However, IM² causes a notable performance drop in fog and snow, likely due to feature redundancy. While MSGF mitigates this to some extent, it remains a concern and will be a key area for future optimization. These results and analyses will be included in the supplementary materials.
@@
MME FAD {IM}2̂ MSGF GFLOPs Inference Speed Params
- - - - 88.19 25ms 18.04M
88.23 (+0.04) 71ms (+46ms) 18.04M (+0M)
91.89 (+3.66) 80ms (+9ms) 20.41M (+2.37M)
232.95 (+141.06) 91ms (+9ms) 55.13M (+34.72M)
255.60 (+22.65) 96ms (+5ms) 61.32M (+6.19M)
Table 11: The ablation experiments on model computation FLOPs, inference speed, and model parameters for each module.
Computational Time and Overhead
We conducted ablation experiments on model computation FLOPs, inference speed, and model parameters for each module. As shown in the table 11, the first row represents the baseline overhead for LiDAR-4DRadar fusion using the naive concatenation of LiDAR and 4DRadar BEV features. Notably, the MME module significantly increases inference time due to computations between the two modal voxels, while the {IM}2 module substantially increases GFLOPs and parameters due to the addition of an extra backbone branch.